OpenStudy (anonymous):

Based on the box-and-whisker plot shown below, match each term with the correct value.

OpenStudy (anonymous):

Median= Range= 25th percentile= 75th percentile= Inter-quartile Range=

OpenStudy (anonymous):

do u understand

OpenStudy (anonymous):

median=12+13/2 =18.5 range=highest-lowest=19-6=13 i only know these...

OpenStudy (anonymous):

i'll try answering the others also :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i got it thanks cutie

OpenStudy (anonymous):

oh, u did? thats great :)

OpenStudy (anonymous):

and 're welcome :)